OpenStudy (anonymous):

Which expression is a CUBE ROOT of -2i? A. cubert(2) (cos(260 degree) + i sin(260 degree)) B. cubert(2) (cos(60 degree) + i sin(60 degree)) C. cubert(2) (cos(90 degree) + i sin(90 degree)) D. cubert(2) (cos(210 degree) + i sin(210 degree))

OpenStudy (anonymous):

Would it be D? (Using improper math)

OpenStudy (anonymous):

To approach this problem, you should write the complex number in polar form.

OpenStudy (anonymous):

Ok! So it would be r = 2?

OpenStudy (anonymous):

For simplicity, let's pull out the -2 for now.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Use Euler's identity to write -i.

OpenStudy (anonymous):

I mean just i.

OpenStudy (anonymous):

Would i = i?

OpenStudy (anonymous):

Remember i is 90 degrees on the complex plane.

OpenStudy (anonymous):

Oh, right!

OpenStudy (anonymous):

So\[i = e^{ i \pi/2} \]

OpenStudy (anonymous):

Now to find the cube root of just i, you can divide the exponent of the polar form by 3.

OpenStudy (anonymous):

\[\sqrt[3]{i} = e^{i \pi/6}\]

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

Now rewrite the polar form into rectangular form.

OpenStudy (anonymous):

And multiply with the cube root of -2, a negative real constant.

OpenStudy (anonymous):

\[\sqrt[3]{2}= e ^{ipi/6}\]Sorry, I don't know how to do this area

OpenStudy (anonymous):

Write it like this: \[e^{i \pi/6} = \cos(\pi/6) + i \sin(\pi/6)\]

OpenStudy (anonymous):

Now in your question, it seems that they are representing the answers in degrees and also they are incorporating the negative factor into the complex part.

OpenStudy (anonymous):

You will need to rewrite the angles in degree form, and also rotate by 90 degrees.

OpenStudy (anonymous):

So \[\frac{ \pi }{ 6} = 30\]

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

And then it would be in the 3rd quadent?

OpenStudy (anonymous):

Oh sorry, I meant rotate by 180 degrees.

OpenStudy (anonymous):

Oh ok! So that makes it 210 degrees, which is answer D!

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Great! Thank you for your help and taking the time to explain this to me!

OpenStudy (anonymous):

It's my pleasure. Good luck.